Good day. i would like to know how to solve this kind of digit problems.
thanks.
The sum of the reciprocals of two numbers is 7. the larger reciprocal exceeds the smaller one by 7/3. find the numbers.

Answer by stanbon(29602)

The sum of the reciprocals of two numbers is 7. the larger reciprocal exceeds the smaller one by 7/3. find the numbers.
------------------------
Equations:
1/x + 1/y = 7
1/x = 1/y + 7/3
-----------------------
1/x = 7-(1/y)
Substitute:
7-(1/y) = (1/y)+(7/3)
(2/y) = 7-(7/3)
2/y = 14/3
y/2 = 3/14
y = 6/14
y = 3/7
----------------
So 1/y = 7/3
---
Substitute to solve for "x":
1/x + (7/3) = 7
1/x = 14/3
x = 3/14
====================
Cheers,
Stan H.